# The Faiss Library

### v2

@@ 4.4 Vector compression benchmark
Figure 4 shows the tradeoff between code size and accuracy for many variants of the codecs. Additive quantizers are the best options for small code sizes. For larger code sizes it is beneficial to independently encode several sub-vectors with product-additive quantizers. LSQ is more accurate than RQ for small codes, but does not scale well to longer codes. Note that product quantizers are a bit less accurate than the additive quantizers but given their low encoding time they remains an attractive option. The scalar quantizers perform well for very long codes and are even faster. The 2-level PQ options are what an IVFPQ index uses as encoding: a first-level coarse quantizer and a second level refinement of the residual (more about this in Section 5.1).
@@ 4.5 Binary indexes
Binary quantization with symmetric distance computations is a pattern that has been commonly used [Wang et al., 2015, Cao et al., 2017]. In this setup, distances are computed in the compressed domain as Hamming distances. (11) reduces to:
$n=\underset{i=1..N}{\mathrm{argmin}}\|C(q)-C_{i}\|$ (17)
where $C(q),C_{i}\in\{0,1\}^{d}$. Although binary quantizers are crude approximations for continuous domain distances, the Hamming distances are integers in $\{0..d\}$, that are fast to compute, do not require any specific context, and are easy to calibrate in practice.
The Faiss IndexBinary indexes support addition and search directly from binary vectors. They offer a compact representation and leverage optimized instructions for distance computations.
@@
  • IndexBinaryHash uses prefix vectors as hashes to cluster the database (rather than spheroids as with inverted lists), and searches only the clusters with closests prefixes.
Finally, a convenience IndexBinaryFromFloat index is provided that simply wraps an arbitrary index and offers a binary vector interface for its operations.
@@ 5 Non-exhaustive search
An early method to do this is Locality Sensitive Hashing (LSH). It amounts to projecting the vectors on a random direction [Datar et al., 2004]. The offsets on that direction are then discretized into buckets where the database vectors are stored. At search time, the buckets nearest to the query vector’s projection are visited. In practice, several projection directions are needed to make it accurate, at the cost of search time. A fundamental drawback of this method is that it is not data-adaptive, although some improvements are possible  [Paulevé et al., 2010].
An alternative way of pruning the search space is to use tree-based indexing. In that case, the dataset is stored in the leaves of a tree [Muja and Lowe, 2014]. When querying a vector, the search starts at the root node. At each internal node, the search descends into one of the child nodes depending on a decision rule. The decision rule depends on how the tree was built: for a KD-tree it is the position w.r.t. a hyperplane, for a hierarchical k-means, it is the proximity to a centroid.
Both in the case of LSH and tree-based methods, the hope is to extend classical database search structures to vector search, because they have a favorable complexity (constant or logarithmic in $N$). However, it turns out that these methods do not scale well for dimensions above 10.
Faiss implements two non-exhaustive search approaches that operate at different memory vs. speed tradeoffs: inverted file and graph-based.
@@ 5.1 Inverted files
IVF indexing is a technique that clusters the database vectors at indexing time. This clustering uses a vector quantizer (the coarse quantizer) that outputs ${K_{\mathrm{IVF}}}$ distinct indices (Faiss’s nlist parameter). The coarse quantizer’s ${K_{\mathrm{IVF}}}$ reproduction values are called centroids. The vectors of each cluster (possibly compressed) are stored contiguously into inverted lists, forming an inverted file (IVF). At search time, only a subset of ${P_{\mathrm{IVF}}}$ clusters are visited (a.k.a. nprobe). The subset is formed by searching the ${P_{\mathrm{IVF}}}$ nearest centroids, as in (2).
@@
The imbalance factor is the relative variance of inverted list sizes [Tavenard et al., 2011]. At search time, if the inverted lists are perfectly balanced (i.e. all have the same length), this factor is 1. If they are unbalanced, the expected number of distance computations is multiplied by this factor.
Figure 5 shows the optimal settings of ${K_{\mathrm{IVF}}}$ for various database sizes. For a small ${K_{\mathrm{IVF}}}=4096$, the coarse quantization runtime is negligible and the search time increases linearly with the database size. For larger datasets it is beneficial to increase the ${K_{\mathrm{IVF}}}$. As in (18), the ratio ${K_{\mathrm{IVF}}}/\sqrt{N}$ is roughly 15 to 20. Note that this ratio depends on the data distribution and the target accuracy. Interestingly, in a regime where ${K_{\mathrm{IVF}}}$ is larger than the optimal setting for $N$ (e.g. ${K_{\mathrm{IVF}}}=2^{18}$ and $N=$5M), the ${P_{\mathrm{IVF}}}$ needed to reach the target accuracy decreases with the dataset size, and so does the search time. This is because when ${K_{\mathrm{IVF}}}$ is fixed and $N$ increases, for a given query vector, the nearest database vector is either the same or a new one that is closer, so it is more likely to be found in a quantization cluster nearer to the query.
With a faster non-exhaustive coarse quantizer (e.g. HNSW) it is even more useful to increase ${K_{\mathrm{IVF}}}$ for larger databases, as the coarse quantization becomes relatively cheap. At the limit, when ${K_{\mathrm{IVF}}}=N$, then all the work is done by the coarse quantizer. However, in that case the limiting factor becomes the memory overhead of the coarse quantizer.
@@
Figure 5: Search time as a function of the database size $N$ for BigANN1M with different ${K_{\mathrm{IVF}}}$ settings. The ${P_{\mathrm{IVF}}}$ is set so that the 1-recall@1 is 90%. The full lines indicate that the coarse quantizer is exact, the dashed lines rely on a HNSW coarse quantizer. For some setting we indicate the ratio ${K_{\mathrm{IVF}}}/\sqrt{N}$
@@
In general, it is more accurate to compress the residuals of the database vectors w.r.t. the centroids [Jégou et al., 2010, Eq. (28)]. This is because the norm of the residuals is lower than that of the original vectors, or because residual encoding is a way to take into account a-priori information from the coarse quantizer. In Faiss, this is controlled via the IndexIVF.by_residual flag, which is set to true by default.
@@
Efficient indexing for maximum inner product search (MIPS) faces multiple issues: the distribution of query vectors is often different from the database vector distribution, most notably in recommendation systems [Paterek, 2007]; the MIPS datasets are diverse, an algorithm that obtains a good performance on some dataset will perform badly on another. Besides, [Morozov and Babenko, 2018] show that using the preprocessing formulas in Section 3 is a suboptimal way of indexing for MIPS.
Several specialized clustering and indexing methods were developed for MIPS [Guo et al., 2020, Morozov and Babenko, 2018]. Instead, Faiss implements a simple modification of k-means clustering, spherical k-means [Dhillon and Modha, 2001], that normalizes the IVF centroids at each iteration. One of the MIPS issues is due to database vectors of very different norms (when they are normalized, MIPS is equivalent to L2 search). High-norm centroids “attract” the database vectors in their clusters, which increases the imbalance factor. Spherical k-means avoids this.
@@
Figure 7: Precision vs. speed tradeoff for the MIPS contriever1M dataset. The compared settings are whether the coarse quantizer assignement is done using L2 distance (default) or IP assignment and whether the k-means clustering does a normalization at each iteration (spherical, IP and L2 assignment are equivalent in that case). The imbalance factors are indicated for each setting.
@@ 5.2 Graph based
Graph-based indexing consists in building a directed graph whose nodes are the vectors to index. At search time, the graph is explored by following the edges towards the nodes that are closest to the query vector. In practice, the search is not greedy but maintains a priority queue with the most promising edges to explore. Thus, the tradeoff at search time is given by the number of exploration steps: higher is more accurate but slower.
A graph-based algorithm is a general framework. Tree-based search or IVF can be seen as special cases of graphs. Graphs can be seen as a way to precompute neighbors for the database vectors, then match the query to one of the vertices and follow the neighbors from there. However, they can also be built to handle out-of-distribution queries [Jaiswal et al., 2022].
Given the search algorithm, taking a pure k-nearest neighbor graph is not optimal because neighbors are too redundant. Therefore, the graph building heuristic consists in balancing edges to nearest neighbors and edges that reach more distant nodes. Most graph methods fix the number of outgoing edges per node, which adjusts the tradeoff between search speed and memory usage. The memory usage per vector breaks down into (1) the possibly compressed vector and (2) the outgoing edges for that vector [Douze et al., 2018].
Faiss implements two graph-based algorithms: HNSW and NSG, in the IndexHNSW and IndexNSG classes, respectively.
@@
The hierarchical navigable small world graph [Malkov and Yashunin, 2018] is an elegant search structure where some vertices (nodes), selected randomly, are promoted to be hubs that are explored first. One of the attractive properties of HNSW is that vectors can be added to it on-the-fly.
@@
The Navigating Spreading-out Graph [Fu et al., 2017] is built from a k-nearest neighbor graph that must be provided on input. At building time, some short-range edges are replaced with longer-range edges. The input k-nn graph can be built with a brute force algorithm or with a specialized method such as NN-descent [Dong et al., 2011] (NNDescent). Unlike HNSW, NSG does not rely on multi-layer graph structures, but uses long connections to achieve fast navigation. In addition, NSG starts from a fixed center point when searching.
Figure 8: Comparison of graph-based indexing methods HNSW (full lines) and NSG (dashes) to index Deep1M. We sweep the tradeoffs between speed and accuracy by varying the number of graph traversal steps (indicated for some of the curves).
@@ Discussion.
Figure 8 compares the speed vs. accuracy for the NSG and HNSW indexes. The main build-time hyper-parameter of the two methods is the number of edges per node, so we tested several settings (for HNSW this is the number of edges on the base level of the hierachical graph). The main search-time parameter is the number of graph traversal steps during search (parameter efSearch for HNSW and search_L for NSG), which we vary to plot each curve. Increasing the number of edges improves the results only to some extent: beyond 64 edges it degrades. NSG obtains better tradeoffs in general, at the cost of a longer build time. Building the input k-NN graph with NN-descent for 1M vectors takes 37 s, and about the same time with brute force search on a GPU (but in that case the k-NN graph is exact). The NSG graph is frozen after the first batch of vectors is added, there is no easy way to add more vectors afterwards.
@@ 5.3 How to choose an index
For most cases choosing an appropriate index can be done by following a process delineated in Figure 9. First, one has to decide whether indexing is needed at all, because in some cases a direct brute force search is the best option. Otherwise, the choice is between IVF and graph-based indexes.
@@
An IVF index can be seen as a special case of graph-based index, especially if a small graph-based index is used as coarse quantizer. Graph-based indices are a good option for indexes where there is no constraint on memory usage, typically for indexes below 1M vectors. Beyond 10M vectors, the construction time typically becomes the limiting factor. For larger indexes, where compression is required to even fit the database vectors in memory, IVF indexes are the only option.
@@ 5.4 Benchmarking indexes
No encoding PQ encoding scalar quantizer
Flat IndexFlat IndexPQ IndexScalarQuantizer
IVF IndexIVFFlat IndexIVFPQ IndexIVFScalarQuantizer
HNSW IndexHSNWFlat IndexHNSWPQ IndexHNSWScalarQuantizer
@@
Beyond a certain scale, search time is determined by the number of distance computations between the query vector and database vectors.
@@
Figure 9: Decision tree to choose a Faiss index. This is for the common case of Euclidean k-nearest neighbor search on CPU. The resulting indexes (in bold) are defined by their factory string (see Section 7.2).
@@ 6 Database operations
In all the experiments above, the indexes are built in one go with all the vectors, while search operations are performed with one batch containing all query vectors. In real settings, the index evolves over time, vectors may be dynamically added or removed, searches may have to take into account metadata, etc. In this section we show how Faiss supports these operations to some extent, mainly on IVF indexes. Specific APIs are available to interface with external storage (Section 7.4) if a finer grained control is required.
@@ 6.1 Identifier-based operations
Faiss indexes support two types of identifiers: sequential ids are based on the order of additions in the index. On the other hand, the user can provide arbitrary 63-bit integer ids along with each vector. The corresponding addition methods for the index are add and add_with_ids. In addition, Faiss supports removing vectors (remove_ids) and updating vector them (update_vectors) by passing the corresponding ids.
Unlike e.g. Usearch [Vardanian, 2022], Faiss does not store arbitrary metadata with the vectors, only 63-bit integer ids can be used (the sign bit is reserved for invalid results).
@@
Sequential indexes (IndexFlatCodes) store vectors as a flat array. They support only sequential ids. When arbitrary ids are needed, the index can be embedded in a IndexIDMap, that translates sequence numbers to arbitrary ids using a int64 array. This enables add_with_ids and returns the arbitrary ids at search time. To also perform id-based operations, a more powerful wrapper class, IndexIDMap2, uses a hash table that maps arbitrary ids back to the sequential ids.
Since graph indexes rely on an embedded IndexFlat Codes to store the actual vectors, they should also be wrapped with the mapping classes. Note however that HNSW does not support suppression and mutation, and that NSG does not even support adding vectors incrementally. Supporting this requires heuristics to re-build the graph when it is mutated, which are implemented in HNSWlib and [Singh et al., 2021] but could be suboptimal indexing-wise.
@@
The IVF indexing structure does supports user-provided ids natively at addition and search time. However, id-based access may require a sequential scan, since the entries are stored in an arbitrary order in the inverted lists. Therefore, the IVF index can optionally maintain a DirectMap, that maps user-visible ids to the inverted list and the offset they are stored in. The map can be an array, which is appropriate for sequential ids, or a hash table, for arbitrary 63-bit ids. Obviously, the direct map incurs a memory overhead and an add-time computation overhead, therefore it is disabled by default. When the direct map is enabled, lookup, removal and update are supported.

### v3

@@ 4.4 Vector compression benchmark
Figure 3 shows the trade-off between code size and accuracy for many variants of the codecs. Additive quantizers are the best options for small code sizes. For larger code sizes, it is beneficial to independently encode several sub-vectors with product-additive quantizers. LSQ is more accurate than RQ for small codes, but does not scale well to longer codes. Note that product quantizers are a bit less accurate than the additive quantizers but given their low encoding time they remains an attractive option. The scalar quantizers perform well for very long codes and are even faster. The 2-level PQ options are what an IVFPQ index uses as encoding: a first-level coarse quantizer and a second level refinement of the residual (more about this in Section 5.1).
@@ 4.5 Binary indexes
Binary quantization with symmetric distance computations is a pattern that has been commonly used [90, 13]. In this setup, distances are computed in the compressed domain as Hamming distances. (11) reduces to:
$n=\underset{i=1..N}{\mathrm{argmin}}\|C(q)-C_{i}\|.$ (17)
where $C(q),C_{i}\in\{0,1\}^{d}$. Hamming distances are integers in $\{0..d\}$. Although they are crude approximations for continuous domain distances, they are fast to compute, do not require any specific context, and are easy to calibrate in practice.
The IndexBinary indexes support addition and search directly from binary vectors. They offer a compact representation and leverage optimized instructions for distance computations.
@@
  • IndexBinaryHash uses prefix vectors as hashes to cluster the database (rather than spheroids as with inverted lists), and searches only the clusters with closest prefixes.
Finally, theIndexBinaryFromFloat is provided for convenience. It wraps an arbitrary index and offers a binary vector interface for its operations.
@@ 5 Non-exhaustive search
A method to do this is Locality Sensitive Hashing (LSH). It amounts to projecting the vectors on a random direction [22]. The offsets on that direction are then discretized into buckets where the database vectors are stored. At search time, only the nearest buckets to the query vector’s projection are visited. In practice, several projection directions are needed to make it accurate, at the cost of search time and memory usage. A fundamental drawback of this method is that it is not data-adaptive, although some improvements are possible [66].
An alternative way of pruning the search space is to use tree-based indexing. In that case, the dataset is stored in the leaves of a tree [62]. When querying a vector, the search starts at the root node. At each internal node, the search descends into one of the child nodes depending on a decision rule. The decision rule depends on how the tree was built: for a KD-tree it is the position w.r.t. a hyperplane, for a hierarchical k-means, it is the proximity to a centroid.
LSH and tree-based methods both aim to extend classical database search structures to vector search, because they have a favorable complexity (constant or logarithmic in $N$). However, these methods do not scale well for dimensions above 10.
Faiss implements two non-exhaustive search approaches that operate at different memory vs. speed trade-offs: inverted file and graph-based.
@@ 5.1 Inverted files
IVF indexing is a technique that clusters the database vectors at indexing time. This clustering uses a vector quantizer (the coarse quantizer) that outputs ${K_{\mathrm{IVF}}}$ distinct indices (the nlist field of the IndexIVF object). The coarse quantizer’s ${K_{\mathrm{IVF}}}$ reproduction values are called centroids. The vectors of each cluster (possibly compressed) are stored contiguously into inverted lists, forming an inverted file (IVF). At search time, only a subset of ${P_{\mathrm{IVF}}}$ clusters are visited (a.k.a. nprobe). The subset is formed by searching the ${P_{\mathrm{IVF}}}$ nearest centroids, as in (2).
@@
The imbalance factor is the relative variance of inverted list sizes [84]. At search time, if the inverted lists all have the same length, this factor is 1. If they are unbalanced, the expected number of distance computations is multiplied by this factor.
Figure 4 shows the optimal settings of ${K_{\mathrm{IVF}}}$ for various database sizes. For a small ${K_{\mathrm{IVF}}}=4096$, the coarse quantization runtime is negligible and the search time increases linearly with the database size. On larger datasets, it is beneficial to increase ${K_{\mathrm{IVF}}}$. As in (18), the ratio ${K_{\mathrm{IVF}}}/\sqrt{N}$ is roughly 15 to 20. Note that this ratio depends on the data distribution and the target accuracy. Interestingly, in a regime where ${K_{\mathrm{IVF}}}$ is larger than the optimal setting for $N$ (e.g. ${K_{\mathrm{IVF}}}=2^{18}$ and $N=$5M), the ${P_{\mathrm{IVF}}}$ needed to reach the target accuracy decreases with the dataset size, and so does the search time. This is because when ${K_{\mathrm{IVF}}}$ is fixed and $N$ increases, for a given query vector, the nearest database vector is either the same or a new one that is closer, so it is more likely to be found in a quantization cluster nearer to the query.
With a faster non-exhaustive coarse quantizer (e.g. HNSW) it is even more useful to increase ${K_{\mathrm{IVF}}}$ for larger databases, as the coarse quantization becomes relatively cheap. At the limit, when ${K_{\mathrm{IVF}}}=N$, then all the work is done by the coarse quantizer. In this scenario, the limiting factor becomes the memory overhead of the coarse quantizer.
@@
Figure 4: Search time as a function of the database size $N$ for BigANN1B with different ${K_{\mathrm{IVF}}}$ settings. The ${P_{\mathrm{IVF}}}$ is set so that the 1-recall@1 is 90%. The full lines indicate that the coarse quantizer is exact, the dashed lines rely on a HNSW coarse quantizer. For some setting we indicate the ratio ${K_{\mathrm{IVF}}}/\sqrt{N}$
@@
In general, it is more accurate to compress the residuals of the database vectors w.r.t. the centroids [44, Eq. (28)]. This is either because the norm of the residuals is lower than that of the original vectors, or because residual encoding is a way to take into account a-priori information from the coarse quantizer. In Faiss, this is controlled via the IndexIVF.by_residual flag, which is set to true by default.
@@
Efficient indexing for maximum inner product search (MIPS) faces multiple issues: the distribution of query vectors is often different from the database vector distribution, most notably in recommendation systems [65]; the MIPS datasets are diverse, an algorithm that obtains a good performance on some dataset will perform badly on another. Besides, [60] show that using the preprocessing formulas in Section 3 is a suboptimal way of indexing for MIPS.
Several specialized clustering and indexing methods were developed for MIPS [36, 60]. Instead, Faiss implements a modification of k-means clustering, spherical k-means [25], which normalizes the IVF centroids at each iteration. One of the MIPS issues is due to database vectors of very different norms (when they are normalized, MIPS is equivalent to L2 search). High-norm centroids “attract” the database vectors in their clusters, which increases the imbalance factor. Spherical k-means is designed to avoid this issue.
@@
Figure 6: Precision vs. speed trade-off for the MIPS contriever1M dataset. The compared settings are whether the coarse quantizer assignement is done using L2 distance (default) or IP assignment and whether the k-means clustering does a normalization at each iteration (spherical, IP and L2 assignment are equivalent in that case). The imbalance factors are indicated for each setting.
@@ 5.2 Graph based
Graph-based indexing consists in building a directed graph whose nodes are the vectors to index. At search time, the graph is explored by following the edges towards the nodes that are closest to the query vector. In practice, the search is not greedy but maintains a priority queue with the most promising edges to explore. Thus, the trade-off at search time is given by the number of exploration steps: higher is more accurate but slower.
A graph-based algorithm is a general framework that can encompass many variants. In particular, tree-based search or IVF can be seen as special cases of graphs. One can see graphs as a way to precompute neighbors for the database vectors, then match the query to one of the vertices and follow the neighbors from there. However, they can also be built to handle out-of-distribution queries [42, 17].
Given this search algorithm, relying on a pure k-nearest neighbor graph is not optimal because neighbors are redundant. Therefore, the graph building heuristic consists in balancing edges to nearest neighbors and edges that reach more distant nodes. Most graph methods fix the number of outgoing edges per node, which adjusts the trade-off between search speed and memory usage. The memory usage per vector breaks down into (1) the possibly compressed vector and (2) the outgoing edges for that vector [29].
Faiss implements two graph-based algorithms: HNSW and NSG, respectively in the IndexHNSW and IndexNSG classes.
@@
The hierarchical navigable small world graph [55] is a search structure where some randomly selected vertices are promoted to be hubs that are explored first. A notable advantage of HNSW is its ability to add vectors on-the-fly.
@@
The Navigating Spreading-out Graph [32] is built from a k-nearest neighbor graph that must be provided on input. At building time, some short-range edges are replaced with longer-range edges. The input k-nn graph can be built with a brute force algorithm or with a specialized method such as NN-descent [26] (NNDescent). Unlike HNSW, NSG does not rely on multi-layer graph structures, but uses long connections to achieve fast navigation. In addition, NSG starts from a fixed center point when searching.
Figure 7: Comparison of graph-based indexing methods HNSW (full lines) and NSG (dashes) to index Deep1M. We sweep the trade-offs between speed and accuracy by varying the number of graph traversal steps (indicated for some of the curves).
@@ Discussion.
Figure 7 compares the speed-accuracy trade-off for the NSG and HNSW indexes. Their main build-time hyperparameter is the number of edges per node, so we tested several settings (for HNSW this is the number of edges on the base level of the hierarchical graph). The main search-time parameter is the number of graph traversal steps during search (parameter efSearch for HNSW and search_L for NSG), which we vary to plot each curve. Increasing the number of edges improves the results until 64 edges, beyond which performance deteriorates. NSG obtains better trade-offs in general, at the cost of a longer build time. Building the k-NN graph with NN-descent for 1M vectors takes 37 s, and about the same time with exact, brute force search on a GPU. The NSG graph is frozen after the first batch of vectors is added, there is no easy way to add more vectors afterwards.
@@ 5.3 How to choose an index
Figure 8: Comparison between the Faiss and SCANN libraries indexing Deep10M in the ann-benchmarks setup (batch mode).
In most cases, choosing an appropriate index can be done by following a process delineated in Appendix A.5. First, one has to decide whether indexing is needed at all: indeed, in some cases a direct brute force search is the best option. Otherwise, the choice is between IVF and graph-based indexes.
@@
Graph-based indices are a good option for indexes where there is no constraint on memory usage, typically for indexes below 1M vectors. Beyond 10M vectors, the construction time typically becomes the limiting factor. For larger indexes, where compression is required to even fit the database vectors in memory, IVF indexes are the only option.
@@ 5.4 Benchmarking indexes
Beyond a certain scale, search time is determined by the number of distance computations performed between the query vector and database vectors.
@@
5.5 Comparison with other libraries
ANN-benchmarks [3] is a codebase that compares several ANNS implementations. We use this setup to compare Faiss with SCANN [36], another industry-standard package for vector search. We run the search in the following setting: batch search on the Deep10M dataset with 10-recall@10 as the metric, where the training is performed on the database vectors. This differs slightly from the evaluation protocol used originally for this dataset.
Figure 8 shows that Faiss is 1.5$\times$ to 4$\times$ faster than SCANN, depending on the operating point. For Faiss we used IVF with a Product Residual Quantizer optimized for SIMD (see Appendix A.3), followed by re-ranking (Section 3.5).
@@ 6 Database operations
In the experiments above, the indexes are built in one go with all the vectors, while search operations are performed with one batch containing all query vectors. In real settings, the index evolves over time, vectors may be dynamically added or removed, searches may have to take into account metadata, etc. In this section we show how Faiss supports some of these operations, mainly on IVF indexes. Specific APIs are available to interface with external storage (Appendix A.4) if fine-grained control is required.
@@ 6.1 Identifier-based operations
Faiss indexes support two types of identifiers: sequential and arbitrary ids. Sequential ids are based on the order of additions in the index. Alternatively, the user can provide arbitrary 63-bit integer ids associated to each vector. The corresponding addition methods for the index are add and add_with_ids. In addition, Faiss supports removing vectors (remove_ids) and updating them (update_vectors) by passing the corresponding ids.
Unlike e.g. Usearch [87], Faiss does not store arbitrary metadata with the vectors, only 63-bit integer ids can be used (the sign bit is reserved for invalid results).
@@
Sequential indexes (IndexFlatCodes) store vectors as a flat array. They support only sequential ids. When arbitrary ids are needed, the index can be embedded in a IndexIDMap, that translates sequence numbers to arbitrary ids using a int64 array. This enables add_with_ids and returns the arbitrary ids at search time. For more id-based operations, a more complete wrapper, IndexIDMap2, maps arbitrary ids back to sequential ids using a hash table.
As graph indexes rely on an embedded IndexFlatCodes to store the actual vectors, they should also be wrapped with the mapping classes. However, HNSW does not support suppression and mutation, and NSG does not even support adding vectors incrementally. Supporting this requires heuristics to rebuild the graph when it is mutated, which are implemented in HNSWlib and [78] but are suboptimal indexing-wise.
@@
The IVF indexing structure supports user-provided ids natively at addition and search time. However, id-based access may require a sequential scan, as the entries are stored in an arbitrary order in the inverted lists. Therefore, the IVF index can optionally maintain a DirectMap, which maps user-visible ids to the inverted list and the offset they are stored in. It supports lookup, removal and update by ids. The map can be an array, which is appropriate for sequential ids, or a hash table, for arbitrary 63-bit ids. Obviously, the direct map incurs a memory overhead and an add-time computation overhead, therefore, it is disabled by default.